5. Consider these properties of chloroform boiling point: 61°C molecular weight: 119.37 g/mole density: 1489 g/ml Cloroform and water are mutually insoluble: they form two layers w hen mixed. wso ml. of water and S0 ml of chloroform are mixed, which layer is on top? Whs? o. We want to convert O.30 moles of phenol into 2.4.6-trinitrophenol using .0 M HNO How many mls of this nitric acid solution should we use? Show your work 7. Consider this reaction CH + 2 CHI → + 21120 +2 Nal - NaOH OH CHj hydroquinone Mol. Wgt. = 110 142 138 Density 2.28 g/mL If you start with 7.0 g hydroquinone and 7.2 mL CHsl, and obtain 5.0 g of p-dimethoxybenzene, what is the percent yield? Show your work.
Answer #1

Q5.

the least dense material will go on top

the density of chloroform > density of water

so

water will be on top

Q6.

0.3 mol of phenol to 2,4,6-trinitrophenol via 3 M of HNO3

find V for itric acid

assuem 100% stoichioemtric ratio

0.3 mol of phenol requires --> 3*0.3 = 0.9 mol of NOx- compounds

so

M = mol/V

V = mol/M = 0.9 / 3

V = 0.3 L = 300 mL required